Ms. Burke invested $15,000 in two accounts, one yielding 5% interest and the other 12%. If she received a total of $1,380 in interest at the end of the year, how much in each account?
Answer by checkley77(12844)

.12X+.05(15,000-X)=1,380
.12X+750-.05X=1,380
.07X=1,380-750
.07X=630
X=630/.07
X=$9,000 IS THE AMOUNT INVESTED @ 12%.
15,000-9,000=6,000 IS THE AMOUNT INVESTED @ 5%
pROOF:
.12*9,000+.05*6,000=1,380
1,080+300=1,380
1,380=1,380